# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 8.2 Subdirect irredundant subpowers
Since $R$ is a counterexample, there are elements $a\neq a^{\prime}$ and tuples $(a_{1},\dotsc,a_{n-1},a),(a_{1},\dotsc$, $a_{n-1},a^{\prime})\in R$. Also, as $R$ is not full, $(c_{1},\dotsc,c_{n})\notin R$ for some tuple in $A^{n}$. Set $T=\{(b_{1},\dotsc,b_{n-1})\mathrel{:}(b_{1},\dotsc,b_{n-1},c_{n})\in R\}$. We show that $T$ together with $I^{\prime}=\{1,2,\dots,n-2\}$ is a smaller counterexample, thus obtaining a contradiction. Indeed, the relation $T$ is proper, as it does not contain $(c_{1},\dotsc,c_{n-1})$. Also, $\operatorname{proj}_{I^{\prime}}(T)$ is the full relation since $\operatorname{proj}_{1,\dotsc,n-2,n}(R)$ is. It remains to show that tuples in $T$ are not determined by their projection to $I^{\prime}$. To this end we consider an auxiliary binary relation $S$ given by $S=\{(a,b)\mathrel{:}(a_{1},\dotsc,a_{n-2},a,b)\in R\}$. This relation is subdirect, as both $\operatorname{proj}_{I}(R)$ and $\operatorname{proj}_{1,\dotsc,n-2,n}(R)$ are full relations. The relation $S$ is also irredundant, because $(a_{n-1},a),(a_{n-1},a^{\prime})\in S$. By the assumptions about $R$, the relation $S$ cannot be proper. Therefore the tuples $(a_{1},\dotsc,a_{n-2},a,c_{n}),(a_{1},\dotsc,a_{n-2},a^{\prime},c_{n})$ are in $R$, implying that $(a_{1},\dotsc,a_{n-2},a),(a_{1},\dotsc,a_{n-2},a^{\prime})$ are in $T$. ∎
@@
Indeed, it is the case if $J\subseteq I$ by the choice of $I$. Suppose for a contradiction that $\operatorname{proj}_{J}(S^{\prime})$ is not full. Then we must have $l\in J$, and, as $\operatorname{proj}_{J\setminus\{l\}}(S^{\prime})$ is the full relation, by Lemma 8.9 every tuple from $S^{\prime}$ is determined by its projection to $J\setminus\{l\}$. This means that $S$ can obtained from $S^{\prime}$ by extending the tuples from $S^{\prime}$ in an arbitrary way, and this is a pp-definition of $S$ from $S^{\prime}$. By the induction hypothesis $S^{\prime}$ is pp-definable from $R_{1},\dotsc,R_{m}$, therefore so is $S$.
@@
We will show that that they are identical. This will imply the result, because, as is easily seen, $S=\operatorname{proj}_{1,\dotsc,l}(U)$, and $U^{\prime}$ is pp-definable from $R_{1},\dotsc,R_{m}$, as $\operatorname{proj}_{1,\dotsc,l-2,l+1}(U)$ is by the induction hypothesis.
It is not hard to see that $U\subseteq U^{\prime}$. Next, note that $U^{\prime\prime}=\operatorname{proj}_{1,\dotsc,l-2,l+1}(U)$ is not full, since $U(c_{1},\dotsc,c_{l-3},a,b,c,d)$ imply that $d=a$. On the other hand, $\operatorname{proj}_{1,\dotsc,l-2}(U)$ is full. Therefore by Lemma 8.9 for any $(a_{1},\dotsc,a_{l},a)\in U$ the value $a$ is determined by $a_{1},\dotsc,a_{l-2}$.
Take a tuple $(a_{1},\dotsc,a_{l},a)\in U^{\prime}$; since $\operatorname{proj}_{1,\dotsc,l-1}(S)$ is the full relation, we have $(a_{1},\dotsc,a_{l-1},c$, $d)\in U$ for some $c,d\in A$. Since $(a_{1},\dotsc,a_{l-2},a)\in\operatorname{proj}_{1,\dotsc,l-2,l+1}(U)$, and in this relation the last value is determined by the first $l-2$ ones, we have $d=a$. Again by Lemma 8.9 the third coordinate of the relation $T$ is determined by the first two coordinates. Therefore, as we have $T(d,a_{l-1},c)$ from the definition of $U$, $T(a,a_{l-1},a_{l})$ from the definition of $U^{\prime}$, and $d=a$, we also obtain $c=a_{l}$. Thus $(a_{1},\dotsc,a_{l},a)\in U$, completing the proof. ∎
As already discussed in Section 4, Theorem 4.7 implies that every algebra $\mathbf{A}$ has at least one of the following properties of its invariant relations.
@@
    $\mathbf{A}$ has no proper irredundant subdirect subpowers.
@@
    $\mathbf{A}$ has a proper irredundant binary subdirect subpower.
@@
    $\mathbf{A}$ has a ternary strongly functional subpower.
Proposition 4.8 shows that $\mathbf{A}$ is abelian in case (3).
@@
Since a tuple of $R$ is determined by any two coordinates, $x_{1}=x_{2}$ implies $v=v^{\prime}$, and this implies $y_{1}=y_{2}$, and vice versa. Since any binary projection of $R$ is full, for any $(x_{1},x_{2})$ we can choose $u$ and then $v$, $v^{\prime}$ such that $(u,v,x_{1}),(u,v^{\prime},x_{2})\in R$. Setting $u^{\prime}=u$, $y_{1}=x_{1}$, and $y_{2}=x_{2}$, we see that $S$ is a reflexive relation on $A^{2}$, in particular, the projection of $S$ to any of the two coordinates is $A^{2}$. Finally, any pair of the form $((x,x),(y,y))$ is in $S$ as witnessed by picking $u$ arbitrarily and then choosing $v=v^{\prime}$ and $u^{\prime}$ appropriately. It follows that the “linkedness” congruence $(S-S)+\dots+(S-S)$ is a congruence on $\mathbf{A}^{2}$ such that one of its classes is the diagonal $\Delta_{A}$, so $\mathbf{A}$ is abelian. ∎
The next theorem produces majority edges in case (1) when $\mu_{\mathbf{A}}$ is not full.
@@
Proof.
We consider the subalgebra $\mathbf{F}$ of $\mathbf{A}^{A^{3}}$ formed by all the ternary term operations of $\mathbf{A}$, i.e., $F=\operatorname{Clo}_{3}(\mathbf{A})$. We set $P=\{(a,b):\operatorname{Sg}_{\mathbf{A}}(a,b)=\mathbf{A}\}$ and define an equivalence relation $\sim$ on $P$ by $(a,b)\sim(a^{\prime},b^{\prime})$ if $t(a,a,b)$ determines $t(a^{\prime},a^{\prime},b^{\prime})$ and vice versa; in other words, if $\{(t(a,a,b),t(a^{\prime},a^{\prime},b^{\prime})):t\in F\}$ is the graph of a bijection (from $A$ to $A$). Note that $\sim$ is indeed an equivalence relation on $P$ and that $(a,b)\sim(a^{\prime},b^{\prime})$ if, and only if, $t(a,b,a)$ and $t(a^{\prime},b^{\prime},a^{\prime})$ determine each other (since coordinates of operations in $F$ can be permuted) and this happens if, and only if, $t(b,a,a)$ and $t(b^{\prime},a^{\prime},a^{\prime})$ determine each other. Let $P^{\prime}$ be a subset of $P$ that contains exactly one representative from each $\sim$-class, let $I=\bigcup_{(a,b)\in P^{\prime}}\{(a,a,b),(a,b,a),(b,a,a)\}$, and finally let $\mathbf{R}\leq\mathbf{A}^{I}$ be the projection of $\mathbf{F}$ onto the set of coordinates $I$.
@@
Since $\mathbf{A}$ has no subdirect proper irredundant subpowers, we get $R=A^{I}$. In particular, there exists $t\in F$ such that $t(a^{\prime},a^{\prime},b^{\prime})=t(a^{\prime},b^{\prime},a^{\prime})=t(b^{\prime},a^{\prime},a^{\prime})=a^{\prime}$ for every $(a^{\prime},b^{\prime})\in P^{\prime}$. It remains to observe that these equalities hold for every $(a,b)\in P$. Indeed, by the definition of $P^{\prime}$ there exists $(a^{\prime},b^{\prime})\in P^{\prime}$ such that $(a,b)\sim(a^{\prime},b^{\prime})$. We have $\operatorname{proj}^{3}_{1}(a^{\prime},a^{\prime},b^{\prime})=a^{\prime}$ and $\operatorname{proj}^{3}_{1}(a,a,b)=a$, therefore $t(a^{\prime},a^{\prime},b^{\prime})=a^{\prime}$ implies $t(a,a,b)=a$. By a similar argument we get $t(a,b,a)=a=t(b,a,a)$ and the proof is concluded. ∎
In case (2) and when $\mathbf{A}$ is simple and $\mu_{\mathbf{A}}$ is not full we get semilattice edges by the following theorem.
@@
Proof.
We assume that $\mu_{\mathbf{A}}$ is not full, otherwise the theorem is vacuously true. We prove the conclusion via a sequence of claims. We start with a basic observation: if $R\leq_{sd}\mathbf{A}^{2}$ and $(a,b),(a^{\prime},b)\in R$ for some $(a,a^{\prime})\notin\mu_{\mathbf{A}}$ then $b$ is in the right center of $R$ (since $\operatorname{Sg}_{\mathbf{A}}(a,a^{\prime})=A$).
@@
Every irredundant, proper $R\leq_{sd}\mathbf{A}^{2}$ is central.
Proof.
Since $\mathbf{A}$ is simple, $R$ needs to be linked.  Therefore, since $\mu_{\mathbf{A}}$ is not the full relation, we have $(a,b),(a^{\prime},b)\in R$ for some $(a,a^{\prime})\notin\mu_{\mathbf{A}}$ which implies that $b$ is in the right center of $R$. The proof for left center is symmetric. ∎
@@
Let $R\leq_{sd}\mathbf{A}^{2}$ be irredundant and proper. If $a$ and $a^{\prime}$ are in the left (right) center of $R$ then $(a,a^{\prime})\in\mu_{\mathbf{A}}$.
Proof.
Suppose not. Then for every $b\in A$ we have $(a,b),(a^{\prime},b)\in R$ which implies that $b$ is in the right center. This cannot happen in a proper $R$. ∎
@@
Let $R\leq_{sd}\mathbf{A}^{2}$ be irredundant and proper. If $a$ is in the left center of $R$ and $a^{\prime}$ is in the right center then $(a,a^{\prime})\in\mu_{\mathbf{A}}$.
Proof.
Suppose, for a contradiction, that $a$ is in the left center of $R$, $a^{\prime}$ is in the right center, and $(a,a^{\prime})\notin\mu_{\mathbf{A}}$. Since $(a,a),(a^{\prime},a^{\prime})\in R$ then, for every $b\in A$, we have $(b,b)\in R$.
Consider any $a^{\prime\prime}$ such that $(a^{\prime\prime},a^{\prime})\notin\mu_{\mathbf{A}}$. Then $a^{\prime\prime}$ is in the left center of $R$, as both $(a^{\prime\prime},a^{\prime\prime})$ and $(a^{\prime\prime},a^{\prime})$ are in $R$. This implies, by the previous claim, that $(a^{\prime\prime},a)\in\mu_{\mathbf{A}}$.
In particular we conclude that $\mu_{\mathbf{A}}$ has two equivalence classes, and that the left center is equal to $a/\mu_{\mathbf{A}}$. By a symmetric argument, the right center is equal to $a^{\prime}/\mu_{\mathbf{A}}$. Altogether, $\mu_{\mathbf{A}}$ is the equivalence with classes $a/\mu_{\mathbf{A}}$ and $a^{\prime}/\mu_{\mathbf{A}}$ which are the left and right centers, respectively.
Also note that no pair $(b,b^{\prime})$ with $b\not\in a/\mu_{\mathbf{A}}$ and $b^{\prime}\not\in a^{\prime}/\mu_{\mathbf{A}}$ is in $R$, as otherwise $(b,b),(b,b^{\prime})\in R$ implies that $b$ is in the left center, which is not the case. Therefore $R=(a/\mu_{\mathbf{A}}\times A)\cup(A\times a^{\prime}/\mu_{\mathbf{A}})$, but then $R\cap-R=\mu_{\mathbf{A}}$ is a proper congruence (recall $|A|>2$) on a simple algebra $\mathbf{A}$, a contradiction. ∎
So far we have shown that each subdirect, irredundant, and proper $R$ is central and both centers are contained in a single $\mu_{\mathbf{A}}$-class $B$. We refer to $B$ as the central $\mu_{\mathbf{A}}$-class of $R$.
@@
Let $R\leq_{sd}\mathbf{A}^{2}$ be irredundant and proper, $B$ its central $\mu_{\mathbf{A}}$-class, and $S\leq_{sd}\mathbf{A}^{2}$ redundant. Then $B+S=B$.
Proof.
The relation $R+S$ is clearly irredundant and proper, its left center is equal to the left center of $R$, and its right center is $C+S$, where $C$ is the right center of $R$. Therefore the central $\mu_{\mathbf{A}}$-class of $R+S$ is $B$ and $C+S\subseteq B$.
The relation $S$ is the graph of an isomorphism from $\mathbf{A}$ to $\mathbf{A}$. Since isomorphism map subuniverses to subuniverses, it maps each $\mu_{\mathbf{A}}$-class to a $\mu_{\mathbf{A}}$-class. In particular $B+S$ is a $\mu_{\mathbf{A}}$-class and since it contains $C+S\subseteq B$, we get $B+S=B$. ∎
@@
Let $R,S\leq_{sd}\mathbf{A}^{2}$ be irredundant and proper and $B,C$ their central $\mu_{\mathbf{A}}$-classes, respectively. Then $B=C$.
Proof.
Suppose, for a contradiction, that $B\neq C$. Let $b,b^{\prime}$ be some elements of the left and right centers of $R$, respectively, and similarly $c,c^{\prime}$ for $S$. Let $T=R\cap S$, and note that $(b,c^{\prime}),(c,b^{\prime})\in T$. As $(b,c)\notin\mu_{\mathbf{A}}$ and $(c^{\prime},b^{\prime})\notin\mu_{\mathbf{A}}$, the relation $T$ is subdirect in $A^{2}$.
Since $B\cap C=\emptyset$ the relation $T$ has no center, and thus by Claim 1 needs to be redundant. The previous claim implies $c^{\prime}\in(B+T)\cap C=B\cap C=\emptyset$, a contradiction. ∎
To finish the proof we take a proper irredundant subdirect binary relation $R$ provided by the assumption and set $B$ to be its central $\mu_{\mathbf{A}}$-class. Take any $b\in B,a\notin B$ and let $S=\operatorname{Sg}_{\mathbf{A}^{2}}((a,b),(b,a))$. The relation $S$ is subdirect (as $(a,b),(b,a)\in S$) and cannot be redundant as $a\in B+S$ would contradict Claim 4. Thus, by Claim 1 and Claim 5, there is $b^{\prime}\in B$ in the left center of $S$. Since $(b^{\prime},b),(a,b)\in S$ we conclude that $b$ is in the right center of $S$, in particular $(b,b)\in S$ and we are done as witnessed by the operation generating $(b,b)$ from the generators $(a,b)$, $(b,a)$. ∎
@@
Proof.
Let us first assume that $\mathbf{A}$ is simple. If case (d) does not apply, then there exists a proper irredundant subdirect subpower of $\mathbf{A}$. By Theorem 4.7, such a relation pp-defines a strongly functional ternary relation or a proper irredundant subdirect binary relation $R$. The former situation leads to case (c) via Proposition 4.8. In the latter case, the binary relation $R\leq_{sd}\mathbf{A}^{2}$ is linked (as $\mathbf{A}$ is simple) and thus it defines a proper left central relation $S\leq_{sd}\mathbf{A}^{2}$ by Proposition 4.1. Now either $\mathbf{A}$ has a nontrivial 2-absorbing subuniverse and we are in case (a) or $S$ witnesses that we are in case (b).
In the general case, take a maximal congruence $\alpha$ of $\mathbf{A}$, apply what we have already proved to the simple algebra $\mathbf{A}/\alpha$, and observe that 2-absorbing subuniverse and centers can be lifted: if $R$ is a 2-absorbing (central) subuniverse of $\mathbf{A}/\alpha$, then the preimage of $R$ under the quotient map $A\to A/\alpha$ is a 2-absorbing (central) subuniverse of $\mathbf{A}$. ∎
@@
Proof.
We prove the claim by induction on the the size of $\mathbf{A}$. If $\mathbf{A}$ has two elements the result follows from the classification of Boolean clones by Post [55], so we assume $|A|\geq 3$.
Suppose that $\mu_{\mathbf{A}}$ is full, that is, each pair of elements in $A$ is connected by proper subuniverses. Since all the proper subalgebras of $\mathbf{A}$ have connected directed graphs of edges by the induction hypothesis, and these edges are trivially also edges in $\mathbf{A}$, it follows that the digraph of edges of $\mathbf{A}$ is connected. Suppose further that $\mu_{\mathbf{A}}$ is not full.
If $\mathbf{A}$ is simple, then each equivalence class of $\mu_{\mathbf{A}}$ is connected by edges by the induction hypothesis. We apply Theorem 4.7 together with Proposition 4.8, Theorem 4.10, or Theorem 4.11 to conclude that either $\mathbf{A}$ is abelian and every pair is an abelian edge, or every pair $(a,a^{\prime})\notin\mu_{\mathbf{A}}$ is a majority edge, or that there exists a $\mu_{\mathbf{A}}$-class $B$ such that every pair $a\notin B,b\in B$ forms a semilattice edge $(a,b)$ (and in every case the witnessing congruence is the identity congruence).
If $\mathbf{A}$ is not simple we consider any maximal congruence $\alpha$ on $\mathbf{A}$. By the induction hypothesis the congruence classes of $\alpha$ as well as $\mathbf{A}/\alpha$ have connected directed graphs of edges. But since any edge $(a/\alpha,b/\alpha)$ in $\mathbf{A}/\alpha$ witnessed by $\theta$ on $\operatorname{Sg}_{\mathbf{A}/\alpha}(a/\alpha,b/\alpha)$ gives rise to edge $(a,b)$ in $\mathbf{A}$ witnessed by the lifted congruence $\theta^{\prime}$ on $\operatorname{Sg}_{\mathbf{A}}(a,b)$, the proof is concluded in the nonsimple case as well. ∎

### v6

@@ 8.2 Subdirect irredundant subpowers
Since $R$ is a counterexample, there are elements $a\neq a^{\prime}$ and tuples $(a_{1},\dotsc,a_{n-1},a),(a_{1},\dotsc$, $a_{n-1},a^{\prime})\in R$. Also, as $R$ is not full, $(c_{1},\dotsc,c_{n})\notin R$ for some tuple in $A^{n}$. Set $T=\{(b_{1},\dotsc,b_{n-1})\mathrel{:}(b_{1},\dotsc,b_{n-1},c_{n})\in R\}$. We show that $T$ together with $I^{\prime}=\{1,2,\dots,n-2\}$ is a smaller counterexample, thus obtaining a contradiction. Indeed, the relation $T$ is proper, as it does not contain $(c_{1},\dotsc,c_{n-1})$. Also, $\operatorname{proj}_{I^{\prime}}(T)$ is the full relation since $\operatorname{proj}_{1,\dotsc,n-2,n}(R)$ is. It remains to show that tuples in $T$ are not determined by their projection to $I^{\prime}$. To this end we consider an auxiliary binary relation $S$ given by $S=\{(a,b)\mathrel{:}(a_{1},\dotsc,a_{n-2},a,b)\in R\}$. This relation is subdirect, as both $\operatorname{proj}_{I}(R)$ and $\operatorname{proj}_{1,\dotsc,n-2,n}(R)$ are full relations. The relation $S$ is also irredundant, because $(a_{n-1},a),(a_{n-1},a^{\prime})\in S$. By the assumptions about $R$, the relation $S$ cannot be proper. Therefore, the tuples $(a_{1},\dotsc,a_{n-2},a,c_{n}),(a_{1},\dotsc,a_{n-2},a^{\prime},c_{n})$ are in $R$, implying that $(a_{1},\dotsc,a_{n-2},a),(a_{1},\dotsc,a_{n-2},a^{\prime})$ are in $T$.
@@
Indeed, it is the case if $J\subseteq I$ by the choice of $I$. Suppose for a contradiction that $\operatorname{proj}_{J}(S^{\prime})$ is not full. Then we must have $l\in J$, and, as $\operatorname{proj}_{J\setminus\{l\}}(S^{\prime})$ is the full relation, by Lemma 8.20 every tuple from $S^{\prime}$ is determined by its projection to $J\setminus\{l\}$. This means that $S$ can be obtained from $S^{\prime}$ by extending the tuples from $S^{\prime}$ in an arbitrary way, and this is a pp-definition of $S$ from $S^{\prime}$. By the induction hypothesis $S^{\prime}$ is pp-definable from $R_{1},\dotsc,R_{m}$, therefore so is $S$.
@@
We will show that they are identical. This will imply the result, because, as is easily seen, $S=\operatorname{proj}_{1,\dotsc,l}(U)$, and $U^{\prime}$ is pp-definable from $R_{1},\dotsc,R_{m}$, as $\operatorname{proj}_{1,\dotsc,l-2,l+1}(U)$ is by the induction hypothesis.
It is not hard to see that $U\subseteq U^{\prime}$. Next, note that $U^{\prime\prime}=\operatorname{proj}_{1,\dotsc,l-2,l+1}(U)$ is not full, since $U(c_{1},\dotsc,c_{l-3},a,b,c,d)$ imply that $d=a$. On the other hand, $\operatorname{proj}_{1,\dotsc,l-2}(U)$ is full. Therefore, by Lemma 8.20 for any $(a_{1},\dotsc,a_{l},a)\in U$ the value $a$ is determined by $a_{1},\dotsc,a_{l-2}$.
Take a tuple $(a_{1},\dotsc,a_{l},a)\in U^{\prime}$; since $\operatorname{proj}_{1,\dotsc,l-1}(S)$ is the full relation, we have that $(a_{1},\dotsc,a_{l-1},c$, $d)\in U$ for some $c,d\in A$. Since $(a_{1},\dotsc,a_{l-2},a)\in\operatorname{proj}_{1,\dotsc,l-2,l+1}(U)$, and in this relation the last value is determined by the first $l-2$ ones, we have $d=a$. Again by Lemma 8.20 the third coordinate of the relation $T$ is determined by the first two coordinates. Therefore, as we have $T(d,a_{l-1},c)$ from the definition of $U$, $T(a,a_{l-1},a_{l})$ from the definition of $U^{\prime}$, and $d=a$, we also obtain $c=a_{l}$. Thus $(a_{1},\dotsc,a_{l},a)\in U$, completing the proof.
As already discussed in Section 4, Theorem 4.7 implies that every algebra has at least one of the following properties of its invariant relations.
@@
    has no proper irredundant subdirect subpowers.
@@
    has a proper irredundant binary subdirect subpower.
@@
    has a ternary strongly functional subpower.
Proposition 4.8 shows that is abelian in case (3).
@@
Since a tuple of $R$ is determined by any two coordinates, $x_{1}=x_{2}$ implies $v=v^{\prime}$, and this implies $y_{1}=y_{2}$, and vice versa. Since any binary projection of $R$ is full, for any $(x_{1},x_{2})$ we can choose $u$ and then $v$, $v^{\prime}$ such that $(u,v,x_{1}),(u,v^{\prime},x_{2})\in R$. Setting $u^{\prime}=u$, $y_{1}=x_{1}$, and $y_{2}=x_{2}$, we see that $S$ is a reflexive relation on $A^{2}$, in particular, the projection of $S$ to any of the two coordinates is $A^{2}$. Finally, any pair of the form $((x,x),(y,y))$ is in $S$ as witnessed by picking $u$ arbitrarily and then choosing $v=v^{\prime}$ and $u^{\prime}$ appropriately. It follows that the “linkedness” congruence $(S-S)+\dots+(S-S)$ is a congruence on such that one of its classes is the diagonal $\Delta_{A}$, so is abelian.
The next theorem produces majority edges in case (1) when $\mu$ is not full.
@@
Proof 8.24.
We consider the subalgebra of formed by all the ternary term operations of , i.e., $F=\operatorname{Clo}_{3}($. We set $P=\{(a,b):\operatorname{Sg}(a,b)=$ and define an equivalence relation $\sim$ on $P$ by $(a,b)\sim(a^{\prime},b^{\prime})$ if $t(a,a,b)$ determines $t(a^{\prime},a^{\prime},b^{\prime})$ and vice versa; in other words, if $\{(t(a,a,b),t(a^{\prime},a^{\prime},b^{\prime})):t\in F\}$ is the graph of a bijection (from $A$ to $A$). Note that $\sim$ is indeed an equivalence relation on $P$ and that $(a,b)\sim(a^{\prime},b^{\prime})$ if, and only if, $t(a,b,a)$ and $t(a^{\prime},b^{\prime},a^{\prime})$ determine each other (since coordinates of operations in $F$ can be permuted) and this happens if, and only if, $t(b,a,a)$ and $t(b^{\prime},a^{\prime},a^{\prime})$ determine each other. Let $P^{\prime}$ be a subset of $P$ that contains exactly one representative from each $\sim$-class, let $I=\bigcup_{(a,b)\in P^{\prime}}\{(a,a,b),(a,b,a),(b,a,a)\}$, and finally let be the projection of onto the set of coordinates $I$.
@@
Since has no subdirect proper irredundant subpowers, we get $R=A^{I}$. In particular, there exists $t\in F$ such that $t(a^{\prime},a^{\prime},b^{\prime})=t(a^{\prime},b^{\prime},a^{\prime})=t(b^{\prime},a^{\prime},a^{\prime})=a^{\prime}$ for every $(a^{\prime},b^{\prime})\in P^{\prime}$. It remains to observe that these equalities hold for every $(a,b)\in P$. Indeed, by the definition of $P^{\prime}$ there exists $(a^{\prime},b^{\prime})\in P^{\prime}$ such that $(a,b)\sim(a^{\prime},b^{\prime})$. We have $\operatorname{proj}^{3}_{1}(a^{\prime},a^{\prime},b^{\prime})=a^{\prime}$ and $\operatorname{proj}^{3}_{1}(a,a,b)=a$, therefore $t(a^{\prime},a^{\prime},b^{\prime})=a^{\prime}$ implies $t(a,a,b)=a$. By a similar argument we get $t(a,b,a)=a=t(b,a,a)$ and the proof is concluded.
In case (2) and when is simple and $\mu$ is not full we get semilattice edges by the following theorem.
@@
Proof 8.25.
We assume that $\mu$ is not full, otherwise the theorem is vacuously true. We prove the conclusion via a sequence of claims. We start with a basic observation: if $R\leq_{sd}$ and $(a,b),(a^{\prime},b)\in R$ for some $(a,a^{\prime})\notin\mu$ then $b$ is in the right center of $R$ (since $\operatorname{Sg}(a,a^{\prime})=A$).
@@
Every irredundant, proper $R\leq_{sd}$ is central.
Proof 8.26.
Since is simple, $R$ needs to be linked.  Therefore, since $\mu$ is not the full relation, we have $(a,b),(a^{\prime},b)\in R$ for some $(a,a^{\prime})\notin\mu$ which implies that $b$ is in the right center of $R$. The proof for left center is symmetric.
@@
Let $R\leq_{sd}$ be irredundant and proper. If $a$ and $a^{\prime}$ are in the left (right) center of $R$ then $(a,a^{\prime})\in\mu$.
Proof 8.27.
Suppose not. Then for every $b\in A$ we have $(a,b),(a^{\prime},b)\in R$ which implies that $b$ is in the right center. This cannot happen in a proper $R$.
@@
Let $R\leq_{sd}$ be irredundant and proper. If $a$ is in the left center of $R$ and $a^{\prime}$ is in the right center then $(a,a^{\prime})\in\mu$.
Proof 8.28.
Suppose, for a contradiction, that $a$ is in the left center of $R$, $a^{\prime}$ is in the right center, and $(a,a^{\prime})\notin\mu$. Since $(a,a),(a^{\prime},a^{\prime})\in R$ then, for every $b\in A$, we have $(b,b)\in R$.
Consider any $a^{\prime\prime}$ such that $(a^{\prime\prime},a^{\prime})\notin\mu$. Then $a^{\prime\prime}$ is in the left center of $R$, as both $(a^{\prime\prime},a^{\prime\prime})$ and $(a^{\prime\prime},a^{\prime})$ are in $R$. This implies, by the previous claim, that $(a^{\prime\prime},a)\in\mu$.
In particular we conclude that $\mu$ has two equivalence classes, and that the left center is equal to $a/\mu$. By a symmetric argument, the right center is equal to $a^{\prime}/\mu$. Altogether, $\mu$ is the equivalence with classes $a/\mu$ and $a^{\prime}/\mu$ which are the left and right centers, respectively.
Also note that no pair $(b,b^{\prime})$ with $b\not\in a/\mu$ and $b^{\prime}\not\in a^{\prime}/\mu$ is in $R$: otherwise $(b,b),(b,b^{\prime})\in R$ implies that $b$ is in the left center, which is not the case. Therefore $R=(a/\mu\times A)\cup(A\times a^{\prime}/\mu)$, but then $R\cap-R=\mu$ is a proper congruence (recall $|A|>2$) on a simple algebra , a contradiction.
So far we have shown that each subdirect, irredundant, and proper $R$ is central and both centers are contained in a single $\mu$-class $B$. We refer to $B$ as the central $\mu$-class of $R$.
@@
Let $R\leq_{sd}$ be irredundant and proper, $B$ its central $\mu$-class, and $S\leq_{sd}$ redundant. Then $B+S=B$.
Proof 8.29.
The relation $R+S$ is clearly irredundant and proper, its left center is equal to the left center of $R$, and its right center is $C+S$, where $C$ is the right center of $R$. Therefore the central $\mu$-class of $R+S$ is $B$ and $C+S\subseteq B$.
The relation $S$ is the graph of an isomorphism from to . Since isomorphism map subuniverses to subuniverses, it maps each $\mu$-class to a $\mu$-class. In particular $B+S$ is a $\mu$-class and since it contains $C+S\subseteq B$, we get $B+S=B$.
@@
Let $R,S\leq_{sd}$ be irredundant and proper and $B,C$ their central $\mu$-classes, respectively. Then $B=C$.
Proof 8.30.
Suppose, for a contradiction, that $B\neq C$. Let $b,b^{\prime}$ be some elements of the left and right centers of $R$, respectively, and similarly $c,c^{\prime}$ for $S$. Let $T=R\cap S$, and note that $(b,c^{\prime}),(c,b^{\prime})\in T$. As $(b,c)\notin\mu$ and $(c^{\prime},b^{\prime})\notin\mu$, the relation $T$ is subdirect in $A^{2}$.
Since $B\cap C=\emptyset$ the relation $T$ has no center, and thus by Claim 1 needs to be redundant. The previous claim implies $c^{\prime}\in(B+T)\cap C=B\cap C=\emptyset$, a contradiction.
To finish the proof we take a proper irredundant subdirect binary relation $R$ provided by the assumption and set $B$ to be its central $\mu$-class. Take any $b\in B,a\notin B$ and let $S=\operatorname{Sg}((a,b),(b,a))$. The relation $S$ is subdirect (as $(a,b),(b,a)\in S$) and cannot be redundant as $a\in B+S$ would contradict Claim 4. Thus, by Claim 1 and Claim 5, there is $b^{\prime}\in B$ in the left center of $S$. Since $(b^{\prime},b),(a,b)\in S$ we conclude that $b$ is in the right center of $S$, in particular $(b,b)\in S$ and we are done as witnessed by the operation generating $(b,b)$ from the generators $(a,b)$, $(b,a)$.
@@
Proof 8.31.
Let us first assume that is simple. If case (d) does not apply, then there exists a proper irredundant subdirect subpower of . By Theorem 4.7, such a relation pp-defines a strongly functional ternary relation or a proper irredundant subdirect binary relation $R$. The former situation leads to case (c) via Proposition 4.8. In the latter case, the binary relation $R\leq_{sd}$ is linked (as is simple) and thus it defines a proper left central relation $S\leq_{sd}$ by Proposition 4.1. Now either has a nontrivial 2-absorbing subuniverse and we are in case (a) or $S$ witnesses that we are in case (b).
In the general case, take a maximal congruence $\alpha$ of , apply what we have already proved to the simple algebra , and observe that 2-absorbing subuniverse and centers can be lifted: if $R$ is a 2-absorbing (central) subuniverse of , then the preimage of $R$ under the quotient map $A\to A/\alpha$ is a 2-absorbing (central) subuniverse of .
@@
Proof 8.32.
We prove the claim by induction on the size of . If has two elements the result follows from the classification of Boolean clones by Post [post41:clones], so we assume $|A|\geq 3$.
Suppose that $\mu$ is full, that is, each pair of elements in $A$ is connected by proper subuniverses. Since all the proper subalgebras of have connected directed graphs of edges by the induction hypothesis, and these edges are trivially also edges in , it follows that the digraph of edges of is connected. Suppose further that $\mu$ is not full.
If is simple, then each equivalence class of $\mu$ is connected by edges by the induction hypothesis. We apply Theorem 4.7 together with Proposition 4.8, Theorem 4.10, or Theorem 4.11 to conclude that either is abelian and every pair is an abelian edge, or every pair $(a,a^{\prime})\notin\mu$ is a majority edge, or that there exists a $\mu$-class $B$ such that every pair $a\notin B,b\in B$ forms a semilattice edge $(a,b)$ (and in every case the witnessing congruence is the identity congruence).
If is not simple we consider any maximal congruence $\alpha$ on . By the induction hypothesis the congruence classes of $\alpha$ as well as have connected directed graphs of edges. But since any edge $(a/\alpha,b/\alpha)$ in witnessed by $\theta$ on $\operatorname{Sg}(a/\alpha,b/\alpha)$ gives rise to edge $(a,b)$ in witnessed by the lifted congruence $\theta^{\prime}$ on $\operatorname{Sg}(a,b)$, the proof is concluded in the nonsimple case as well.